OpenStudy (anonymous):

The length of a rectangle is 8 inches greater than the width. The perimeter is 80 inches. What is the width of the rectangle? Lily decided to let w stand for the width of the rectangle. Which problem-solving step was this? A. Check your answer and present the solution. B. Come to an answer. C. Understand the problem. D. Gather your resources.

OpenStudy (unklerhaukus):

\[ P_R=2\times( w +l)\]

OpenStudy (unklerhaukus):

\[l=8+w\]

OpenStudy (carniel):

80 = 8+W

OpenStudy (unklerhaukus):

\[80=2\times(w+8+w)\]